Is there a point that lies on all three equations? What is it?
y=-x+9
y=4x
y=(2)/(5)x

Answer by Alan3354(69443)

Is there a point that lies on all three equations? What is it?
y=-x+9
y=4x
y=(2)/(5)x
--------------
Find the intersection of the 1st 2 eqns.
y=-x+9
y=4x
4x = x-9
x = -3, y = -12
----------------
y = (2/5)x does not.
y = 2/(5x) does not.